OpenStudy (akashdeepdeb):

\(mn = 2010020020010002 = 2^1×7^5×11^5×13^5\) How many possible (m,n) ordered pairs are there?

Parth (parthkohli):

I don't know how counting works, but it's definitely to do something with separating the list of factors into two different sets.

OpenStudy (akashdeepdeb):

I think it is just a Combinatorics Question.

OpenStudy (akashdeepdeb):

@ganeshie8

Parth (parthkohli):

Yes it is. For example if you have to do \(mn = 20\) then factorize \(2 \times 2 \times 5\). Find the number of ways to separate the set \(\{2,2,5\}\) into two individual unique sets.

ganeshie8 (ganeshie8):

\(mn = 2010020020010002 = 2^1×7^5×11^5×13^5\) number of factors = \((1+1)(5+1)(5+1)(5+1)\)

ganeshie8 (ganeshie8):

you can find the possible ordered pairs ^

OpenStudy (akashdeepdeb):

How did you calculate that? And is, number of factors = possible ordered pairs?

Parth (parthkohli):

{2,2} and {5} {2,5} and {2} also include {2,2,5}, {1} {1}, {2,2,5}

OpenStudy (akashdeepdeb):

@ParthKohli That was for 20 How do i do it for: 2010020020010002 ?

Parth (parthkohli):

generalize a method by looking at smaller numbers. I did a similar problem but that was long ago. with that, I leave you. you're in safe hands though (ganesh!)

ganeshie8 (ganeshie8):

Oh wait a sec you're right ! total number of ordered pairs = total number of factors

OpenStudy (akashdeepdeb):

Ganesh, Why is it (432)^2. And how did you get 432 ??

OpenStudy (akashdeepdeb):

(1+1)(5+1)(5+1)(5+1) = 432

ganeshie8 (ganeshie8):

\((1+1)(5+1)(5+1)(5+1)\) should be the final answer

OpenStudy (akashdeepdeb):

Ok, how did you get (1+1)(5+1)(5+1)(5+1) ? T.T

ganeshie8 (ganeshie8):

\(mn = 2010020020010002 = 2^1×7^5×11^5×13^5\)

OpenStudy (akashdeepdeb):

Yes.

ganeshie8 (ganeshie8):

\(m\) can hav 2^0 or 2^1 in its prime factorization : (1+1) ways

ganeshie8 (ganeshie8):

\(m\) can hav 7^0 or 7^1 or 7^2 or 7^3 or 7^4 or 7^5 in its prime factorization : (5+1) ways

ganeshie8 (ganeshie8):

\(\cdots\)

OpenStudy (akashdeepdeb):

What do you actually mean by that?

OpenStudy (akashdeepdeb):

(5+1) ways ?

OpenStudy (anonymous):

wat if mn= (17^3*11)(...) ??

ganeshie8 (ganeshie8):

yes the prime power for 7, can occur in (5+1) = 6 ways in m

OpenStudy (akashdeepdeb):

Okay. So why is N(m,n) = 432 ?

ganeshie8 (ganeshie8):

In general : if \(N = p_1^{a_1}p_2^{a_2}\cdots p_ra^r\) is the prime factorization of \(N\) then total number of factors = \((a_1+1)(a_2+1)\cdots (a_r+1)\)

OpenStudy (anonymous):

right ! forgot about that

ganeshie8 (ganeshie8):

choosing \(m\), uniquely determines the other factor\(n\), so total number of ordered pairs equals total number of factors

OpenStudy (akashdeepdeb):

How do I actually understand that formula? Any link? And I understand how No. of ordered pairs = No. of factors! :D

ganeshie8 (ganeshie8):

let me try explaining u quick before giving the link

ganeshie8 (ganeshie8):

\(N = p_1^{a_1}p_2^{a_2}\cdots p_r^{a_r} \)

ganeshie8 (ganeshie8):

First observation : \(p_1, ~p_2, .... p_r\) are the factors of \(N\) as they divide \(N\) evenly

OpenStudy (akashdeepdeb):

Yes.

ganeshie8 (ganeshie8):

Next, think at finding how many ways the `power` of a specific prime factor can occur in a factor.

ganeshie8 (ganeshie8):

in our present problem, 7^5 is there in factorization of N so, 7^0 or 7^1 or 7^2 or 7^2 or 7^3 or 7^4 or 7^5 are also factors of \(N\)

OpenStudy (akashdeepdeb):

For 7^5 1 * mn 7 * x 49 * y 343 * z 7^4 * (some number) 7^5 * (2*13^5*11^5) For a specific prime factor.

OpenStudy (akashdeepdeb):

Okay. Yes. Then?

ganeshie8 (ganeshie8):

yes a power of 7 can occur in 6 ways

ganeshie8 (ganeshie8):

next look at other primes

ganeshie8 (ganeshie8):

p^r can occur in (r+1) ways

ganeshie8 (ganeshie8):

multiply them all to get all the prime power combinations

OpenStudy (anonymous):

its like this :- u have 20 number , first u wanna chose one # =m or 2 # =m or 3 num =m or 4#=5 ...... or 20#=m :3

OpenStudy (akashdeepdeb):

Yes. All the prime factors can occur in (r+1) ways. What about their combinations now?

ganeshie8 (ganeshie8):

thats the fundamental principal of counting

ganeshie8 (ganeshie8):

if 7^5 can be chosen in 6 ways andd 11^5 can be chosen in 6 ways then you can choose both of them in 6*6 ways

OpenStudy (akashdeepdeb):

Oh I think I get it now! "This is when the concept clicks". *__* Thank you very much @ganeshie8 !

ganeshie8 (ganeshie8):

\(a_1 + 1\) ways for \(p_1^{a_1}\) \(a_2 + 1\) ways for \(p_2^{a_2}\) \(\cdots \) \(a_r + 1\) ways for \(p_r^{a_r}\) so total combinations = \((a_1+1)(a_2+1)\cdots (a_r+1)\)

OpenStudy (anonymous):

@ AkashdeepDeb i couldn't correct early it should be (1+1)(5+1)(5+1)(5+1)/2 = 216

OpenStudy (anonymous):

For e.g 20 =2^2*5^1 So the total number of factors are (2+1)(1+1)=6 (1,2,4,5,10,20) but the number of ways in which 20 can be expressed is 6/2 = 3 hence we have (1,20) ,(4,5) and (2,10)

OpenStudy (phi):

I think ganeshie explained it well, but here is what I thought: they want the number of ordered pairs which means (for example) (1,2) is different from (2,1), and all we have to count are the number of unique m's we can form m can either use or not use the factor of 2: use either 1 or 2. That is 2 choices m can use from 0 up to 5 sevens, so multiply by 6 (number of choices) same for 11 and 13 we get 2*6*6*6= 432

OpenStudy (anonymous):

here i don't think so to consider (1,2) and (2,1) to be different pairs...

OpenStudy (phi):

The question states: How many possible (m,n) ordered pairs are there? *ordered* pair means the order of m and n matter, i.e. (m,n) is different from (n,m)